### v1

# Resolving Task Objective Conflicts in Unified Multimodal Understanding and Generation via Task-Aware Mixture-of-Experts

Abstract
Unified multimodal large language models (MLLMs) based on end-to-end autoregressive (AR) transformers effectively integrate both understanding and generation tasks within a single framework. However, intrinsic Task Objective Conflicts between high-level semantic abstraction in understanding and fine-grained detail preservation in generation pose significant challenges, often leading to suboptimal trade-offs and task interference. Existing solutions, such as decoupling shared visual encoders, fall short of fundamentally resolving these conflicts due to inherent AR architecture. In this paper, we propose a novel approach that decouples internal components of AR to resolve task objective conflicts. Specifically, we design UTAMoE, a Unified Task-Aware Mixture-of-Experts (MoE) framework that decouples internal AR modules via a Task-Aware MoE Layer to create task-specific optimization subpaths. To enhance task differentiation while maintaining overall coordination, we introduce a novel Two-Stage Training Strategy. Extensive experiments on multimodal benchmarks demonstrate that UTAMoE mitigates task objective conflicts, achieving state-of-the-art performance across various tasks. Visualizations and ablation studies further validate the effectiveness of our approach.
Resolving Task Objective Conflicts in Unified Multimodal Understanding and Generation via Task-Aware Mixture-of-Experts
Jiaxing Zhang Sichuan University zjxing9972@gmail.com Xinyi Zeng Sichuan University perperstudy@gmail.com Hao Tang††thanks: Corresponding author. Peking University haotang@pku.edu.cn
1 Introduction
Recently, the emergence of end-to-end autoregressive (AR) Transformer architectures Wu et al. (2024c); Xie et al. (2024) have emerged as an effective solution for unified modeling of multimodal understanding and generation, enabling genuine integration of both tasks within a single framework. However, a fundamental challenge arises from the intrinsic divergence in task objectives: understanding tasks aim to distill high-level semantic representations, whereas generation tasks require preserving and integrating fine-grained details. This task objective conflict often leads to interference during training, making it difficult for models to effectively balance both tasks.
Recent efforts, such as Janus Wu et al. (2024a); Ma et al. (2024); Chen et al. (2025), attempt to mitigate this conflict by decoupling the shared visual encoder, enabling specific visual representations and reducing feature-level interference. However, such strategies do not fundamentally resolve the divergence in task objectives. As illustrated in Figure 1 a), current models struggle to achieve an optimal trade-off between understanding and generation. This is reflected in issues like semantic degradation in comprehension (e.g., answering "What is on the grass?" with a generic label "animal"), or missing fine-grained details in generation (e.g., omitting the "puppy" in "A boy sits on a chair with a puppy beside him").
The encoder-level decoupling proposed by Janus motivates our further exploration of whether internal modules within the autoregressive architecture can be decoupled to better balance conflicting objectives across tasks. In contrast to encoder-level decoupling, internal decoupling requires a careful trade-off between maintaining architectural unification and enabling task-specific specialization. Excessive decoupling may undermine shared contextual understanding and parameter efficiency, while insufficient decoupling may fail to alleviate cross-task interference. Therefore, a mechanism is needed to enable task-specific while maintaining unified training and computational efficiency.
To this end, we identify Mixture of Experts (MoE) Dai et al. (2024); Liu et al. (2024a) as a promising strategy for structural decoupling. By dynamically activating different experts to construct task-specific subpaths, MoE allows disentangling feature spaces and reducing cross-task interference—thereby enabling efficient and specialized processing in unified multimodal architectures.
In this paper, we propose UTAMoE, a Unified Task-Aware Mixture-of-Experts framework designed to address Task Objective Conflicts in multimodal understanding and generation. The key idea is to construct task-oriented optimization subpaths within AR using a Task-Aware MoE Layer. Additionally, we propose a Two-Stage Training Strategy to further enhance task-specific differentiation while maintaining overall performance balance. As shown in Figure 1 b), UTAMoE yields more desirable outcomes in both tasks—accurate recognition (e.g., "puppy") in understanding and improved detail preservation in generation.
We conduct extensive experiments on multiple multimodal benchmarks and observe that UTAMoE achieves substantial improvements in semantic recognition accuracy for understanding tasks, along with enhanced detail preservation for generation tasks. Both quantitative, qualitative, and ablation studies validate the effectiveness of decoupling the internal AR architecture in mitigating task objective conflict. We hope that UTAMoE offer novel perspectives for unified multimodal tasks and offer valuable insights into effectively addressing task objective conflicts identified in this work.
Our main contributions are summarized below:
We propose a novel approach to decouple internal components of autoregressive (AR) architectures, effectively mitigating Task Objective Conflicts.
We present UTAMoE, a Task-Aware MoE framework that enables task-specific optimization within unified multimodal models.
We introduce a Two-Stage Training Strategy that helps UTAMoE achieve state-of-the-art performance across multiple benchmarks.
2 Related Work
Mixture of Experts (MoE).
Recent studies Dai et al. (2024); Liu et al. (2024a); Cai et al. (2024) have shown MoE’s effectiveness when integrated with language models. For instance, Mixtral-MoE 8x7B Jiang et al. (2024) excels in scaling model capacity and improving task performance. DeepSeek introduced Top-K routing and shared experts, enhancing MoE’s task allocation efficiency. Variants like Chen et al. Chen et al. (2023b) enable multitask learning for simple tasks through embedding analysis, while Li et al. Li et al. (2025) assign different experts for different modalities, proving expert specialization’s feasibility. These advancements strongly support MoE’s integration into unified multimodal frameworks.
Multimodal Understanding.
Multimodal large language models (MLLMs), such as LLaVA Liu et al. (2023), MiniGPT-4 Zhu et al. (2023), and InstructBLIP Panagopoulou et al. (2023), are strong in processing multimodal information. They project features from pretrained modality-specific encoders (e.g., CLIP Radford et al. (2021)) into LLMs to enable cross-modal reasoning. However, these models focus mainly on visual perception and lack native generative ability. Recent studies Ge et al. (2023); Jin et al. (2023); Sun et al. (2023) explore extending MLLMs to other domains, which presents an area worth further exploration.
Visual Generation.
Visual generation is a rapidly evolving field where diffusion-based methods have excelled in text-to-image and text-to-video tasks. These methods operate in a continuous latent space obtained via a VAE encoder, predicting Gaussian noise added to the latent representations. Inspired by natural language processing, autoregressive (AR) models use transformer architectures to predict discrete visual tokens Ramesh et al. (2021); Esser et al. (2021). These models discretize visual data and apply GPT-style mechanisms for generation. Additionally, masked prediction models adopt BERT-like strategies to improve generation efficiency by predicting masked input portions. Such approaches have been extended to more complex multimodal tasks, including video generation Yu et al. (2023a).
Unified Multimodal Models.
Unified multimodal understanding and generation models are regarded as highly effective for enabling seamless reasoning and content generation across diverse modalities Team (2024); Zhou et al. (2024). However, existing approaches, whether based on autoregressive (AR) architectures with shared encoders Wu et al. (2024c); Team (2024) or decoupled ones Wu et al. (2024a); Ma et al. (2024); Chen et al. (2025), often suffer from task objective conflicts when jointly addressing both understanding and generation tasks. To address this conflict, our UTAMoE leverages the MoE mechanism to establish task-specific optimization pathways within AR architecture, thereby effectively addressing task objective conflict.
3 The Proposed Method
Recently, Unified multimodal large language models based autoregressive (AR) Transformers Team (2024); Wu et al. (2024c) address understanding and generation tasks in a single framework. However, the conflicts arising from the fundamental divergence between the objectives of understanding and generation tasks reveal the inherent limitations of the autoregressive (AR) architecture. Janus decouples the shared visual encoder to produce task-specific visual representations, effectively mitigating interference at the visual feature level across tasks. However, this approach does not fully resolve the task objective conflict. To address this issue, we propose a novel approach centered on decoupling the internal components of the autoregressive (AR) architecture. Specifically, we design a Task-Aware MoE Layer, as detailed in subsection 3.1, to replace the conventional feed-forward network (FFN) layer, thereby forming our unified architecture—UTAMoE. This module incorporates two key mechanisms: Hierarchical Expert Routing for task-specific expert selection and Hybrid Expert Collaboration for inter-task synergy.
Additionally, to fully leverage unique strengths of different expert modules and improve overall generalization of the framework, we propose a novel two-stage training strategy, detailed in subsection 3.2, to further enhance performance of UTAMoE.
3.1 Task-Aware MoE Layer
To better accommodate task objective conflict of understanding and generation tasks, we introduce a modular design based on a Mixture-of-Experts (MoE) framework. Specifically, the expert are divided into two groups, each tailored to understanding and generation tasks, respectively. To enhance task-specific specialization, we propose a Hierarchical Expert Routing mechanism, which comprises a task-aware router and a conventional dynamic assignment router. This mechanism enables more accurate routing of task-relevant tokens to the appropriate expert. Furthermore, to ensure coherence and synergy across task objectives, we incorporate a shared expert independent of the task-specific groups. As a core part of the Hybrid Expert Collaboration mechanism, the shared expert facilitates cross-task information exchange and helps maintain a balanced representation between understanding and generation tasks.
3.1.1 Hierarchical Expert Routing
To achieve fine-grained expert selection tailored to different task objectives, we introduce Hierarchical Expert Routing, consisting of a Task-Aware Router and a Dynamic-Assignment Router.
Task-Aware Router.
Given an input token representation , we first classify it into one of the two task-specific expert groups via a softmax-based classifier:
This essentially acts as a hard routing mechanism, directly determining the group that the token should be routed to.
Dynamic-Assignment Router.
Once the expert group is determined, we further compute the relevance between the token and each expert in the group, and select the top- experts based on the scores:
Here, denotes the group-specific expert scoring matrix, where is the number of experts in each group.
3.1.2 Hybrid Expert Collaboration
After selecting experts, it processes input in a weighted combination:
To leverage both task-specific and global knowledge, we integrate the output of a shared expert:
where is a learnable hyperparameter that balances the contributions.
Loss Function.
The overall training objective includes both task-specific losses and group loss:
where is the ground-truth for the task group, is the loss for task , and , are hyperparameters controlling the loss weights.Specifically, corresponds to the understanding task, while controls the generation task.
3.2 Training Strategy
To fully leverage the unique capabilities of different expert modules and enhance the generalization ability of the overall framework, we propose a Two-stage Training strategy, as illustrated in Figure 3, to construct UTAMoE. In the following paragraphs, we detail the objectives and provide a comprehensive explanation of each training stage.
Stage 1: Training Task-Specific Experts.
This stage focuses on acquiring task-specific expert through dedicated training on individual tasks. The objective is to obtain expert weights that perform well on their respective tasks. Specifically, we adopt the generative entropy loss as the primary training criterion. During this process, the self-attention layers are frozen, and only the feed-forward networks (FFNs) are updated, resulting in well-specialized task-specific experts.
Stage 2: LoRA-Based Fine-Tuning of UTAMoE.
In this stage, we first replace the conventional FFN layers with the Task-Aware MoE layer. Then, we integrate the previously obtained expert weights and perform joint fine-tuning of UTAMoE using a mixture of understanding and generation tasks. Specifically, the Task-Aware MoE layer utilizes a task-aware router to dynamically select the appropriate task group, within which a Top-K strategy is applied to choose task-specific experts. We then perform end-to-end fine-tuning using the LoRA approach to optimize the entire model while maintaining high parameter efficiency.
Through LoRA fine-tuning, our UTAMoE efficiently adapts to diverse tasks while maintaining high parameter efficiency, thereby enhancing the model’s generalization capability and enabling effective training in multi-task settings.
4 Experiments
In this section, we introduce comprehensive experiments to evaluate our method.Firstly, we outline our experimental setup, including model architecture and training configurations. Next, we evaluate our model on various understanding and generation benchmarks. We also provide qualitative results to further illustrate the strengths of our approach. Finally, we analyze the expert utilization for both task-specific expert and shared expert.
4.1 Experimental Setups
Model Architecture.
We adopt JanusPro as the base model, whose language component supports a maximum sequence length of 4096. For multimodal understanding, we use SigLIP-Large-Patch16-384 Zhai et al. (2023) as the image encoder. For visual generation, the encoder utilizes a codebook of size 16,384 and downsamples inputs by a factor of 16. Both understanding and generation adapters are implemented as two-layer MLPs. Additionally, we integrate the Any-Resolution technique based Liu et al. (2024d) to support inputs at varying resolutions.
Training Configurations.
Since our training data are subsets of those used in Wu et al. (2024a) and we use JanusPro as base model, the vision tower weights are kept frozen throughout all training stages, as they have already been sufficiently pretrained and generalize well across similar data distributions. For training task-specific experts, we utilize ShareGPT4V Chen et al. (2024b) dataset for understanding tasks and MidJourney Vivym (2023) dataset for generation tasks. During the LoRA-based fine-tuning stage, a mixed dataset combining both sources is employed. Regarding the MoE configuration, each task group is assigned two experts, and a Top-K routing strategy is applied within the group. Additionally, inspired by DeepSeekMoE Dai et al. (2024), one shared expert is included across all groups to encourage generalization. All experiments are conducted on a single NVIDIA V100 GPU (32GB). We set , , and , with the remaining hyperparameters detailed in Table 1.
| Hyperparameter | Stage 1 | Stage 2 |
| Batch size | 2 | 2 |
| Learning rate | ||
| Optimizer | AdamW (, ) | |
| LR Scheduler | Cosine | |
| Weight decay | 0.0 | 0.0 |
| Training steps | 200 | 400 |
| Type | Model | Params | POPE↑ | MME-P↑ | MMB↑ | SEED↑ | GQA↑ | MMMU↑ | MM-Vet↑ |
| Und. Only | LlaVA-v1.5-Phi-1.5 Xie et al. (2024) | 1.3B | 84.1 | 1128.0 | - | - | 56.5 | 307 | - |
| MobileVLM Chu et al. (2023) | 1.4B | 84.5 | 1196.2 | 53.2 | - | 56.1 | - | - | |
| MobileVLM-V2 Chu et al. (2024) | 1.4B | 84.3 | 1302.8 | 57.7 | - | 59.3 | - | - | |
| MobileVLM Chu et al. (2023) | 2.7B | 84.9 | 1288.9 | 59.6 | - | 59.0 | - | - | |
| MobileVLM-V2 Chu et al. (2024) | 2.7B | 84.7 | 1440.5 | 63.2 | - | 61.1 | - | - | |
| LlaVA-Phi Zhu et al. (2024) | 2.7B | 85.0 | 1335.1 | 59.8 | - | - | - | 28.9 | |
| LLaVA Liu et al. (2023) | 7B | 76.3 | 809.6 | 38.7 | 33.5 | - | - | 25.5 | |
| LLaVA-v1.5 Liu et al. (2024c) | 7B | 85.9 | 1510.7 | 64.3 | 58.6 | 62.0 | 35.4 | 31.1 | |
| InstructBLIP Dai et al. (2023) | 7B | - | - | 36.0 | 53.4 | 49.2 | - | 26.2 | |
| Qwen-VL-Chat Bai et al. (2023a) | 7B | - | 1487.5 | 60.6 | 58.2 | 57.5 | - | - | |
| Emu3-Chat Wang et al. (2024b) | 8B | 85.2 | 1244 | 58.5 | 68.2 | 60.3 | 31.6 | 37.2 | |
| InstructBLIP Dai et al. (2023) | 13B | 78.9 | 1212.8 | - | - | 49.5 | - | 25.6 | |
| Und. and Gen. | DreamLLM† Dong et al. (2023) | 7B | - | - | - | - | - | - | 36.6 |
| LaViT† Jin et al. (2023) | 7B | - | - | - | - | 46.8 | - | - | |
| MetaMorph† Tong et al. (2024) | 8B | - | - | 75.2 | 71.8 | - | - | - | |
| NEXT-GPT† Wu et al. (2024b) | 13B | - | - | - | - | - | - | - | |
[…]
| SDHo-512 Xie et al. (2024) | 1.3B | 80.0 | 1097.2 | - | - | 58.0 | 26.7 | - | |
| D-DiT Li et al. (2024c) | 2.0B | 84.0 | 1124.7 | - | - | 59.2 | - | - | |
| Gemini-Nano-1 Team et al. (2023) | 1.8B | - | - | - | - | - | 26.3 | - | |
| TokenFlow-XL Qu et al. (2024) | 13B | 86.8 | 1545.9 | 68.9 | 68.7 | 62.7 | 38.7 | 40.7 | |
| LWM Liu et al. (2024b) | 7B | 75.2 | - | - | 44.8 | - | - | 9.6 | |
| VILA-U Wu et al. (2024c) | 7B | 85.8 | 1401.8 | - | 59.0 | 60.8 | - | 33.5 | |
| Chameleon Team (2024) | 7B | - | - | - | - | - | 22.4 | 8.3 | |
| Janus Wu et al. (2024a) | 1.5B | 87.0 | 1338.0 | 69.4 | 63.7 | 59.1 | 30.5 | 34.3 | |
| JanusPro Chen et al. (2025) | 1.5B | 86.2 | 1444.0 | 75.5 | 68.3 | 59.3 | 36.3 | 39.8 | |
| UTAMoE (Ours) | 3B | 87.2 | 1534.1 | 79.3 | 72.4 | 61.5 | 41.7 | 46.7 |
4.2 Quantitative Evaluation
Multimodal Understanding.
To assess multimodal understanding capabilities, we evaluate our model on widely recognized image-based vision-language benchmarks, which include POPE Li et al. (2023b), MME Fu et al. (2024), MMB Liu et al. (2024e), SEED Li et al. (2024a), GQA Hudson and Manning (2019), MMMU Yue et al. (2024) and MM-Vet Yu et al. (2023b).
Visual Generation.
For evaluating visual generation capabilities, we use GenEval Ghosh et al. (2023). GenEval is a challenging benchmark for text-to-image generation, designed to reflect the comprehensive generative abilities of visual generation models by offering a detailed instance-level analysis of their compositional capabilities.
Comparison Methods.
As shown in Table 2 and Table 3, we categorize the models used for comparison into three groups based on their capabilities. (i) Understanding Only: This group includes models designed solely for multimodal understanding, such as the LLaVA series (e.g., LLaVA Liu et al. (2023), LLaVA-v1.5 Liu et al. (2024c)), the MobileVLM series (e.g., MobileVLM Chu et al. (2023), MobileVLM-V2 Chu et al. (2024)), InstructBLIP Dai et al. (2023), Qwen-VL-Chat Bai et al. (2023a), and Emu3-Chat Wang et al. (2024b). (ii) Generation Only: This group contains models that focus exclusively on visual generation, including LlamaGen Sun et al. (2024), LDM Rombach et al. (2022), SDv1.5 Rombach et al. (2022), PixArt- Chen et al. (2023a), SDv2.1 Rombach et al. (2022), DALL·E 2 Ramesh et al. (2022), Emu3-Gen Wang et al. (2024b), SDXL Podell et al. (2023), DALL·E 3 Betker et al. (2023), and SD3-Medium Esser et al. (2024). (iii) Understanding and Generation: This group comprises models capable of both multimodal understanding and generation, such as DreamLLM Dong et al. (2023), LaViT Jin et al. (2023), MetaMorph Tong et al. (2024), NEXT-GPT Wu et al. (2024b), Show-o-256 Xie et al. (2024), SDHo-512 Xie et al. (2024), D-DiT Li et al. (2024c), Gemini-Nano-1 Team et al. (2023), ILLUME Wang et al. (2024a), TokenFlow-XL Qu et al. (2024), LWM Liu et al. (2024b), VILA-U Wu et al. (2024c), Chameleon Team (2024), and the Janus-series models Wu et al. (2024a); Chen et al. (2025); Ma et al. (2024).
| Type | Method | Single Obj. | Two Obj. | Counting | Colors | Position | Color Attri. | Overall |
| Gen. Only | LlamaGen Sun et al. (2024) | 0.71 | 0.34 | 0.21 | 0.58 | 0.07 | 0.04 | 0.32 |
| LDM Rombach et al. (2022) | 0.92 | 0.29 | 0.23 | 0.70 | 0.02 | 0.05 | 0.37 | |
| SDv1.5 Rombach et al. (2022) | 0.97 | 0.38 | 0.35 | 0.76 | 0.04 | 0.06 | 0.43 | |
| PixArt- Chen et al. (2023a) | 0.98 | 0.50 | 0.44 | 0.80 | 0.08 | 0.07 | 0.48 | |
| SDv2.1 Rombach et al. (2022) | 0.98 | 0.51 | 0.44 | 0.85 | 0.07 | 0.17 | 0.50 | |
| DALL-E2 Ramesh et al. (2022) | 0.94 | 0.66 | 0.49 | 0.77 | 0.10 | 0.19 | 0.52 | |
| Emu3-Gen Wang et al. (2024b) | 0.98 | 0.71 | 0.34 | 0.81 | 0.17 | 0.21 | 0.54 | |
| SDXL Podell et al. (2023) | 0.98 | 0.74 | 0.39 | 0.85 | 0.15 | 0.23 | 0.55 | |
| DALL-E 3 Betker et al. (2023) | 0.96 | 0.87 | 0.47 | 0.83 | 0.43 | 0.45 | 0.67 | |
| SD3-Medium Esser et al. (2024) | 0.99 | 0.94 | 0.72 | 0.89 | 0.33 | 0.60 | 0.74 | |
| Und. and Gen. | SEED-X† Li et al. (2024a) | 0.97 | 0.58 | 0.26 | 0.80 | 0.19 | 0.14 | 0.49 |
[…]
| D-DiT Li et al. (2024c) | 0.97 | 0.80 | 0.54 | 0.76 | 0.32 | 0.50 | 0.65 | |
| LWM Liu et al. (2024b) | 0.93 | 0.41 | 0.46 | 0.79 | 0.09 | 0.15 | 0.47 | |
| Transfusion Zhou et al. (2024) | - | - | - | - | - | - | 0.63 | |
| TokenFlow-XL Qu et al. (2024) | 0.95 | 0.60 | 0.41 | 0.81 | 0.16 | 0.24 | 0.55 | |
| Chameleon Team (2024) | - | - | - | - | - | - | 0.39 | |
| Janus Wu et al. (2024a) | 0.97 | 0.68 | 0.30 | 0.84 | 0.46 | 0.42 | 0.61 | |
| JanusPro Chen et al. (2025) | 0.98 | 0.82 | 0.51 | 0.89 | 0.65 | 0.56 | 0.73 | |
| UTAMoE (Ours) | 0.99 | 0.84 | 0.55 | 0.91 | 0.62 | 0.65 | 0.76 |
All in all, our UTAMoE, based on Chen et al. (2025), achieves state-of-the-art performance in multimodal understanding tasks, and delivers strong results in visual generation. This success is primarily attributed to our redesigned Mixture-of-Experts (MoE) mechanism, which effectively mitigates the task objective conflicts between multimodal understanding and generation, thereby significantly enhancing model’s overall capabilities.
4.3 Qualitative Evaluation
Figure 5 presents qualitative results for both multimodal understanding (left) and visual generation (right). UTAMoE demonstrates outstanding comprehension capabilities when processing inputs from diverse contexts, showcasing its strong capacity for multimodal semantic modeling, particularly in Basic Visual Description, Visual Question Answering (VQA), and Text Recognition tasks. The right side presents representative visual generation results across four categories: basic spatial relationships, natural landscapes, human characters, and imaginative scenes. Although the output resolution of UTAMoE is limited to 384 × 384, the model effectively captures the semantic intent of prompts, generating coherent, well-structured, and detail-rich images. These results highlight the superiority of our approach: UTAMoE not only models high-level semantic information in multimodal understanding tasks but also faithfully preserves fine-grained visual details in generation tasks, thereby mitigating inter-task objective conflicts.
4.4 Expert Load Analysis
Figure 4 illustrates expert utilization in understanding task, visualized by averaging values from randomly selected samples across multiple experimental runs. As shown in 4, Expert-0 and Expert-1 belong to the understanding group, with their combined activation rate exceeding 60% in most cases. Experts 2 and 3, belonging to the generation group, exhibit low activation rates across most layers, indicating that our task-aware router effectively routes tokens to the intended expert groups.Even when occasional routing errors increase generation group activations, the shared expert’s influence can partially compensate for meeting the task requirements. This further validates the effectiveness of our routing mechanism and shared expert design.
5 Ablation Study
In this section, we conduct a comprehensive ablation study to investigate the contributions of key components in our proposed framework. We focus on analyzing the effectiveness of Task-Aware MoE and the impact of our training strategy.
5.1 Effectiveness of Task-Aware MoE
| Model | Task-Aware Router | Shared | Understanding | Generation | ||
| + Group Loss | Expert | MMMU | POPE | Gen.eval | ||
| A | ✗ | ✗ | 36.8 | 85.4 | 0.70 | |
| B | ✓ | ✗ | 39.9 | 86.9 | 0.72 | |
| C | ✓ | ✓ | 41.7 | 87.2 | 0.76 | |
| Model | MoE | Train. Method | Expert Init. | Epochs to Convergence | MMMU | POPE | Gen.eval |
| D | ✓ | Single Stage | pure | 94 | 36.5 | 85.3 | 0.68 |
| E | ✓ | Two Stage | from Stage1 | 76 | 41.7 | 87.2 | 0.76 |
To assess the effectiveness of each component in the Task-Aware MoE, we progressively construct a series of ablation models, as shown in Table 5. We begin with Model A as the baseline, which excludes two critical components: the Task-Aware Router (including Group Loss) and the Shared Expert. Model B introduces the Task-Aware Router, enabling dynamic expert group selection based on task-specific requirements. Building on this, Model C incorporates the Shared Expert module, which promotes cross-task knowledge sharing and enhances the model’s generalization capability.
We evaluate understanding performance using the MMMU and POPE benchmarks, and assess generation quality with the Gen.eval metric, reporting the “overall” score as the primary evaluation criterion. As shown in Table 5, each added component yields consistent performance gains, with Model C achieving the best results across all tasks. This ablation demonstrates the effectiveness and necessity of each module in enhancing both task-specific adaptation and generalization capabilities within the Task-Aware MoE framework.
5.2 Effectiveness of Training Strategy
We investigate the impact of different training strategies on model performance through two distinct model variants presented in Table 5. To ensure a fair comparison, all models share the same architecture and are evaluated on identical task sets. The comparative results between models D and E demonstrate that the two-stage training strategy employed by model E yields significant improvements across various benchmarks. This indicates that incorporating dedicated training stages for handling distinct tasks not only enhances model performance but also accelerates convergence, underscoring the strategic advantage of involving experts in task-specific training.
6 Conclusion
In this paper, we propose a novel approach to the problem of Task Objective Conflicts in unified multimodal autoregressive models. The foundational idea of our approach is to decouple the internal components of autoregressive architecture. Specifically, we design UTAMoE, featuring a Task-Aware Mixture-of-Experts layer that creates task-specific optimization subpaths, effectively disentangling the conflicting objectives of understanding and generation tasks. Furthermore, we develop a Two-Stage Training Strategy that enhances task-specific differentiation while balancing objectives of distinct tasks. Extensive experimental results demonstrate that UTAMoE achieves state-of-the-art performance across various benchmarks.
7 Limitations
The diversity and inconsistency of current datasets, including variations in modality types, annotation quality, and task definitions, pose significant challenges to unified training and fair evaluation. Additionally, reliance on large pretrained backbone models results in substantial computational costs.
Appendix A Experimental Settings and Details
A.1 Details of Model Architecture
In our experiments, UTAMoE adopts a unified architecture for both multimodal understanding and generation tasks, inheriting the overall design from Chen et al. (2025). The architecture consists of decoupled vision encoder, aligners, text encoder, large language model (LLM), text decoder and image generation head. Except for LLM and decoupled vision encoder, all other components remain consistent across tasks.
Vision Encoder and Aligner.
For understanding tasks, we employ a vision encoder based on the CLIPVisionTower architecture with siglip-large-patch16-384 as the backbone, processing images at a resolution of 384×384. For generation tasks, we utilize a VQ-16 encoder with a downsampling factor of 16, producing 16,384 image tokens with a feature dimension of 8. Prior to inputting into the LLM, visual features from both encoders are projected into a unified embedding space via their respective alignment modules. Each alignment module is implemented as a two-layer MLP. the understanding aligner accepts inputs of dimension 1024 and outputs 2048-dimensional embeddings with GELU activation, while the generation aligner maps 8-dimensional inputs into the same 2048-dimensional space. The aligned features from both tasks are subsequently integrated into language model for subsequent multimodal fusion and processing.
Core Language Model.
Our language model is based on DeepSeek-LLM, supporting a maximum sequence length of 4,096 tokens. Building upon this foundation, we incorporate a Task-Aware Mixture-of-Experts (MoE) Layer to replace the original Transformer’s feed-forward network (FFN) modules, thereby resolving task objective conflict. Specifically, the model comprises 24 Transformer layers, each with the FFN replaced by our designed Task-Aware MoE layer containing two group-specific experts and one shared expert. The model has a hidden dimension of 2,048 and an intermediate dimension of 5,632, and is trained using bfloat16 precision. The vocabulary size is 102,400, and the maximum positional embedding supports sequence lengths up to 16,384 tokens. The multi-head attention mechanism is compatible with the LLaMA architecture.
Vision Head.
For the text-to-image generation task, we incorporate a vision head module to transform the language model outputs into discrete image tokens. This module is configured with an embedding dimension of 2048 and generates a total of 16,384 image tokens, enabling high-fidelity image reconstruction and visualization.
A.2 Training Dataset Details
We employ ShareGPT4V to train understanding tasks, while MidJourney is utilized for training generation tasks. Detailed descriptions of both datasets are provided below.
ShareGPT4V is a carefully curated collection of 102,000 high-quality image-text pairs, gathered from diverse sources including LAION, Conceptual Captions (CC), SBU Captions, SAM, COCO, web-landmark, web-celebrity, and Wikiart. Each pair was selected and refined to ensure accuracy, relevance, and quality. The GPT-4 Vision model was used during curation to analyze and improve captions, ensuring they align well with the images and are informative. This dataset is a valuable resource for researchers in computer vision and natural language processing, supporting tasks like image captioning, visual question answering, multimodal learning, content-based image retrieval, and art or landmark recognition. The ShareGPT4V dataset is stored in JSON format (see A.2), with each entry following a VQA format that includes image paths, questions, and answers—making it easy to load and process for model training.
MidJourney provides over 20 million prompt-image pairs encompassing diverse styles, compositions, and model versions, making it a valuable resource for training and research in text-to-image generation models. The dataset includes high-quality images paired with their corresponding prompts, as illustrated in A.2. In this work, we utilize a carefully selected subset of this dataset.
A.3 Evaluation Benchmarks
A.3.1 Multimodal Understanding
POPE: Evaluates object hallucination by converting hallucination detection into binary yes/no questions about object presence in MSCOCO images. It supports both annotated and raw images using automatic segmentation tools. Analysis shows 9% ambiguous and 5.5% incorrect annotations in the dataset.
MME: Covers 14 sub-tasks across perception, reasoning, and language understanding. It features diverse question types, instruction styles, and metrics, providing comprehensive diagnostic insights into MLLM performance.
MMBench: A multiple-choice evaluation benchmark in English and Chinese. It uses a Circular Evaluation strategy—shifting answer positions across multiple queries—to reduce bias and assess model consistency under varied choice orderings.
SEED-Bench: Tests interleaved vision-language comprehension and generation across three difficulty levels. Tasks include image-text understanding and multimodal instruction following, emphasizing compositional reasoning and complex visual input handling.
GQA: Focuses on compositional reasoning in visual question answering. Questions are generated from scene graphs to assess understanding of spatial relationships, attributes, and object roles in real-world images.
MMMU: Contains 11.5K multimodal questions drawn from textbooks and exams across six academic disciplines. Evaluates subject-specific reasoning with both visual and textual input, simulating high-level, expert-domain tasks.
MM-Vet: Assesses 16 combinations of six core skills: object recognition, commonsense knowledge, spatial understanding, text generation, OCR, and visual math. Uses LLM-based metrics for consistent scoring of open-ended answers.
A.3.2 Visual Generation
GenEval: Includes diverse challenges such as single-object rendering, object pairs, counting, spatial relations, and color accuracy. Evaluations are typically conducted using object detection models to assess generation fidelity and alignment with prompts.
DPG-Bench: Features prompts from multiple sources with an average length of 67 words, offering significant compositional and linguistic complexity. Generated images are evaluated via visual question answering models to test semantic alignment and content accuracy.
A.4 Image Processing: Any-resolution Strategy
Current unified models for multimodal understanding and generation typically employ fixed-resolution visual encoders. For instance, JanusPro utilizes the SigLIP Zhai et al. (2023) with a fixed 384384 input resolution for multimodal understanding tasks, which inevitably degrades performance on fine-grained tasks like text recognition due to loss of detailed information. Similarly, in visual generation, such resolution constraints lead to insufficient detail preservation in synthesized images.
Previous efforts Bai et al. (2023b); Li et al. (2023a) to enhance resolution often rely on positional embedding interpolation followed by fine-tuning Vision Transformer (ViT) backbone on large-scale image-text datasets. However, this approach not only demands substantial computational resources but also rigidly fixes the acceptable input resolution during inference. To overcome these challenges, we propose Any-Resolution strategy based Liu et al. (2024d) as illustrated in Figure 6.
Given an input image , we begin by dividing into local image patches.
Then, for each image patch , we randomly select one transformation from a set of predefined geometric transformations (e.g., rotation, shear, etc.) to apply to the patch.
Each transformed image is then encoded by visual encoder to produce visual tokens:
Simultaneously, the resized original image is encoded as:
All encoded features are flattened and concatenated to form a comprehensive representation:
Appendix B Supplementary Experiments
B.1 Validation of Task Objective Conflict
Comparative Analysis.
To investigate potential conflicts between task objectives, we conduct a comparative experiment between single-task and multi-task training. Specifically, we fine-tune two variants based on Chen et al. (2025) : (1) a single-task model, where each task is trained independently with its own objective function; and (2) a multi-task model, where all tasks are jointly optimized using a weighted combination of task-specific losses. We select POPE and Gen.Eval as representative benchmarks to cover both multimodal understanding and visual generation tasks. As shown in Table 6, if multi-task model shows noticeable performance drops on certain tasks and fails to achieve balanced performance across all tasks, this indicates the presence of inherent conflicts between task objectives.
| Task | POPE | GenEval | POPE | Gen.Eval | |
| (a) | Und. | 86.9 | – | – | – |
| (b) | Gen. | – | 0.74 | – | – |
| (c) | Both | 86.2 | 0.73 | -0.7 | -0.01 |
The results indicate the existence of task objective conflict. However, focusing solely on final performance metrics may overlook subtle interactions. To gain deeper insights, we further analyze the dynamics of the loss functions during training. Specifically, we examine the variations in the loss functions used for multimodal understanding and visual generation tasks. This detailed investigation helps confirm and understand the nature of these conflicts beyond just evaluation scores.
Loss Function Analysis.
To further verify the task objective conflict, we analyze the dynamics of the loss functions during multi-task training. Specifically, we track the loss values of multimodal understanding and visual generation tasks over training iterations to identify whether decreases in one task’s loss correspond to increases in another’s. Our understanding task uses cross-entropy loss, while the generation task employs mean squared error (MSE) loss. If significant mutual suppression in loss curves, it indicates the existence of optimization conflicts between tasks.
From Figure 7, we observe a certain degree of antagonism between the two loss functions, which further supports the existence of task conflicts. However, these experiments do not fundamentally verify the conflict, and addressing this at a deeper level remains a goal for our future work.
B.2 Generation Performance on DPG-Bench
DPG-Bench (Dense Prompt Graph Benchmark) comprises 1,065 complex and densely structured prompts, designed for evaluating visual generation capabilities, particularly in terms of fine-grained semantic alignment under linguistically and conceptually challenging scenarios. Results are summarized in Table 7.
| Method | Global | Entity | Attribute | Relation | Other | Overall |
| SDv1.5 Rombach et al. (2022) | 74.63 | 74.23 | 75.39 | 73.49 | 67.81 | 63.18 |
| PixArt- Chen et al. (2023a) | 74.97 | 79.32 | 78.60 | 82.57 | 76.96 | 71.11 |
| SDXL Podell et al. (2023) | 83.27 | 82.43 | 80.91 | 86.76 | 80.41 | 74.65 |
| Hunyuan-DiT Li et al. (2024b) | 84.59 | 80.59 | 88.01 | 74.36 | 86.41 | 78.87 |
| PixArt- Chen et al. (2024a) | 86.89 | 82.89 | 88.94 | 86.59 | 87.68 | 80.54 |
| Emu3-Gen Wang et al. (2024b) | 85.21 | 86.68 | 86.84 | 90.22 | 83.15 | 80.60 |
| DALL-E 3 Betker et al. (2023) | 90.97 | 89.61 | 88.39 | 90.58 | 89.83 | 83.50 |
| SD3-Medium Esser et al. (2024) | 87.90 | 91.01 | 88.83 | 80.70 | 88.68 | 84.08 |
| Janus Wu et al. (2024a) | 82.33 | 87.38 | 87.70 | 85.46 | 86.41 | 79.68 |
| JanusPro Chen et al. (2025) | 87.58 | 88.63 | 88.17 | 88.98 | 88.30 | 82.63 |
| UTAMoE (Ours) | 87.95 | 89.07 | 89.31 | 89.33 | 88.84 | 84.11 |
B.3 Comprehensive Analysis of Expert
The visualization analysis in subsection 4.4 illustrates the expert utilization patterns of our trained UTAMoE on the understanding task. However, several important questions remain to be explored. In the following subsection, we focus on addressing one key issue: What is the optimal ratio between group-specific experts and shared experts?
Optimal Ratio of Group-Specific and Shared Experts.
To explore the impact of the ratio between group-specific experts and shared experts on model performance, we conduct a series of experiments varying the G S ratio. Table 8 summarizes the results on POPE and GenEval benchmarks under different expert configurations. By analyzing these outcomes, we aim to identify the optimal balance that maximizes overall task performance and effectively leverages both specialized and shared knowledge within the model. Notably, we do not adopt the 3:1 ratio proposed by DeepSeek Dai et al. (2024), largely because we have redesigned the routing strategy, which alters the interaction between experts and affects the ideal ratio.
| G:S Ratio | POPE | GenEval | |
| (d) | 1:0 | 85.5 | 0.68 |
| (e) | 0:1 | 86.9 | 0.74 |
| (f) | 1:1 | 87.1 | 0.74 |
| (g) | 2:1 | 87.2 | 0.76 |
| (h) | 3:1 | 86.2 | 0.73 |
As presented in Table 8, our preliminary findings suggest that an optimal ratio of group-specific to shared experts is approximately 2. We emphasize that this conclusion is preliminary, and further extensive investigations will be conducted to validate and refine this observation in future work.
Appendix C Additional Visualizations
C.1 Multimodal Understanding.
In addition to the task types illustrated in Figure 5, several other representative tasks are widely studied in the field of multimodal understanding. Image-Text Matching focuses on determining whether a given image and a textual description semantically correspond to each other. Landmark Recognition focuses on detecting and classifying well-known landmarks or notable locations within an image. Visual Dialog requires engaging in a multi-turn conversation by jointly reasoning over both the visual content and the dialog history to generate coherent and contextually appropriate responses. To further highlight the robustness of our model’s multimodal understanding capabilities, we present additional qualitative examples in Figure 12.
C.2 Visual Generation.
Beyond scenarios presented in Figure 5, text-to-image generation also encompasses a wide range of additional contexts, including urban versus indoor, abstract versus concrete, and human-centered versus object-centered perspectives. The results are shown in Figure 13 below.