### v2

# Resolving Task Objective Conflicts in Unified Model via Task-Aware Mixture-of-Experts

Abstract.
Recently, multimodal understanding (MMU) and text-to-image generation (T2I) have been integrated into a single autoregressive (AR) architecture, achieving initial unification. However, existing works focus on representation-level studies and overlook potential conflicts in AR architectures’ internal information flow during training different tasks. Motivated by this gap, we identify a deeper issue, Task Objective Conflict (TOC), arising from AR architectures’ internal information flow, which causes negative transfer and catastrophic forgetting when training MMU and T2I jointly. To address this issue, we proposed UniDecouple, which decouples internal modules for different tasks to construct task-specific optimization subpaths. To implement UniDecouple, we employ a Task-Aware Mixture-of-Experts (TA-MoE), comprising Hierarchical Expert Routing and Hybrid Expert Collaboration, trained in two stages: first to build task-specific experts, then jointly fine-tuned to balance specialization and overall coordination. Extensive experiments on both understanding and generation benchmarks demonstrate that UniDecouple preserves strong understanding ability while achieving generation quality comparable to state-of-the-art methods, offering a new perspective for unified multimodal modeling.
Key words and phrases:
Unified Model, Autoregressive, Mixture-of-Experts1. Introduction
Human cognition, where understanding and generation intricately interleave, provides the foundation for reasoning and thought Islam and Moushi (2025); Wu et al. (2024a); Tong et al. (2024). Inspired by this mechanism, researchers have sought unified models that integrate multimodal understanding and generation within a single framework. Yet progress remains divided: autoregressive (AR) based models Zhu et al. (2024b); Liu et al. (2024d); Zhu et al. (2023); Bai et al. (2023) drive multimodal understanding (MMU), while diffusion approaches Podell et al. (2023); Wang et al. (2024c); Rombach et al. (2022) dominate text-to-image generation (T2I). To bridge these divergent paradigms, recent works (e.g., DALL·E Reddy et al. (2021)) employ AR based models for T2I, resulting in a preliminarily unified paradigm and prompting further investigations Team (2024); Chern et al. (2024); Dong et al. (2023); Wang et al. (2024b); Wu et al. (2024c) into extending AR frameworks to support both MMU and T2I. Yet, a well-balanced trade-off between enhancing generation quality and retaining understanding ability remains challenging for AR based models.
A widely recognized challenge for AR based unified models arises from representation level Wu et al. (2024a); Ma et al. (2024); Chen et al. (2025). MMU relies on semantic abstraction to capture high-level information, whereas T2I demands fine-grained detail preservation to faithfully render the image. The conflict in representation requirements has been widely recognized, and to reconcile it, early approaches first adopted a single pixel or semantic encoder to learn shared representations Wu et al. (2024c); Lin et al. (2025). Nevertheless, relying on a single encoder often biases the extracted features toward one task and cannot adequately satisfy these distinct objectives. Subsequently, JanusPro Chen et al. (2025) decouples the encoders, leveraging both pixel and semantic encoders to derive task-specific representations, effectively resolving the representation conflict.
Nevertheless, prior work on unifying MMU and T2I within AR based architectures focuses on surface-level representation conflicts, overlooks the deeper issues inherent in the architecture itself. Consequently, we focus on the internal information flow and propose that conflicts also emerge within the architecture Sener and Koltun (2018); Yu et al. (2020). Subsequently, we conduct theoretical analyses and empirical experiments to validate this hypothesis, demonstrating that these conflicts, termed Task Objective Conflict (TOC), indeed exist. Specifically, training MMU and T2I simultaneously using AR based approaches can lead to negative transfer, where learning on one task (e.g., understanding) inadvertently degrades performance on another (e.g., generation), and catastrophic forgetting, where previously acquired task-specific knowledge is partially lost.
To mitigate TOC, we propose UniDecouple. As illustrated in Fig. 1(c), UniDecouple uniquely integrates with the AR paradigm by leveraging a Mixture of Experts (MoE), whose intrinsic task-specific routing mechanism provides dedicated subpaths for each task. Nevertheless, the conventional MoE architecture remains inadequate for well resolving these conflicts, as it lacks explicit mechanisms to orchestrate task-specific experts and reconcile the divergent objectives of MMU and T2I. To this end, we design Task-Aware MoE (TA-MoE), which incorporates key components: Hierarchical Expert Routing, enabling more precise task-specific pathway selection, and Hybrid Expert Collaboration, facilitating the integration of complementary knowledge. We also introduce a two-stage training strategy: training experts for task-specific skills and jointly fine-tuned to balance specialization and overall coordination.
1
We empirically validate the effectiveness of UniDecouple through a series of experiments. In Sec. 4.2, we first validate our hypothesis by analyzing the behavior of the loss function and examining how TOC manifests during training. In Secs. 4.3, we then conduct both quantitative and qualitative evaluations on mainstream understanding and generation benchmarks, demonstrating that UniDecouple consistently achieves strong performance across tasks. Finally, in Sec. 5, ablation studies confirm the contributions of our proposed TA-MoE architecture and the two-stage training strategy. Overall, these results demonstrate the robustness and effectiveness of UniDecouple. Our main contributions are summarized below:
We validate the existence of Task Objective Conflict (TOC) in widely adopted autoregressive (AR) frameworks from both theoretical and empirical perspectives.
We propose UniDecouple, a Task-Aware MoE (TA-MoE) framework that enables task-specific optimization within unified multimodal models.
Our comprehensive experiments show that UniDecouple achieves generation quality on par with current state-of-the-art methods while effectively preserving the model’s understanding ability.
2. Related Work
Mixture of Experts (MoE)
Recent studies Dai et al. (2024); Liu et al. (2024b); Cai et al. (2024) have demonstrated the effectiveness of MoE when integrated with large language models. For instance, Mixtral-MoE 8x7B Jiang et al. (2024) excels in scaling model capacity and improving task performance. DeepSeek introduced Top-K routing and shared experts, enhancing MoE’s task allocation efficiency. Variants such as Chen et al. Chen et al. (2023a) enable multitask learning for simple tasks through embedding analysis, while Li et al. Li et al. (2025) assign different experts for different modalities, proving the feasibility of expert specialization.
Multimodal Understanding (MMU)
Multimodal large language models (MLLMs), such as Flamingo Alayrac et al. (2022), LLaVA Liu et al. (2023), MiniGPT-4 Zhu et al. (2023), BLIP-2 Li et al. (2023b), InstructBLIP Panagopoulou et al. (2023), Qwen2-VL Wang et al. (2024a), and Intern-VL Chen et al. (2024c), demonstrate strong capabilities in processing multimodal information. They typically align features from pretrained modality-specific encoders (e.g., CLIP Radford et al. (2021)) using methods such as Q-Former and connectors, and then map them into the LLM to enable cross-modal reasoning. Recent studies Ge et al. (2023); Jin et al. (2023); Sun et al. (2023) have attempted to extend MLLMs beyond perception to generation and other domains, opening new research directions.
Text-to-Image Generation (T2I)
Generation has undergone a series of methodological advancements. Initially, diffusion-based methods generated high-quality samples by iteratively denoising Gaussian perturbations in continuous latent spaces Ho et al. (2020); Nichol and Dhariwal (2021). The traditional denoising process relied on U-Net, it has gradually evolved to Transformer-based prediction method Peebles and Xie (2023). Diffusion models excel in generation fidelity and stability, but their multi-step iterative process results in slow inference. To improve efficiency, flow matching Lipman et al. (2022) generative models were introduced, learning continuous probability flows to more directly approximate the data distribution while retaining the advantages of diffusion. Subsequently, autoregressive (AR) models began to tackle visual generation tasks Reddy et al. (2021), discretizing image into token sequences and employing Transformer to predict each token sequentially. These advancements lay a solid foundation for the development of subsequent unified multimodal models.
Unified Multimodal Models.
Unified multimodal understanding and generation models, based on AR, are regarded as highly effective for enabling seamless reasoning and content generation across diverse modalities Team (2024); Zhou et al. (2024). However, existing approaches, whether based on autoregressive (AR) architectures with shared encoders Wu et al. (2024c); Team (2024) or decoupled ones Wu et al. (2024a); Ma et al. (2024); Chen et al. (2025), often suffer from task objective conflicts (TOC) when jointly addressing both understanding and generation tasks. These conflicts can lead to negative transfer and also slow down convergence during joint training. To address this conflict, our UniDecouple leverages the MoE mechanism to establish task-specific optimization pathways within AR architecture, thereby effectively addressing TOC.
2
3. UniDecouple
In this section, we first provide a theoretical analysis of the Task Objective Conflicts (TOC) in Section 3.1, which highlight the inherent limitations of unified autoregressive (AR) frameworks when addressing both MMU and T2I tasks. Building on this foundation, we introduce the fundamental component of our proposed Unidecouple architecture: Task-Aware Mixture of Experts (TA-MoE) Layer, which decouples the internal structure to enable task-specific expert selection and effective cross-task collaboration (details presented in Section 3.2). Furthermore, to fully exploit the complementary strengths of different expert modules and enhance the overall generalization capability of the framework, we propose a two-stage training strategy in Section 3.3.
3.1. Task Obejective Conflict
We formalize the TOC within the context of AR modeling, highlighting how sequential dependency and attention masking amplify gradient conflict between tasks.
Consider an AR model parameterized by , trained on a fixed dataset to perform two tasks: MMU and T2I. The model predicts a token sequence under an auto-regressive factorization:
where is modeled by causal attention. The two tasks differ in objective scope: MMU abstracts semantics from observed tokens, while T2I reconstructs unobserved tokens over longer horizons, with task losses defined as:
MMU mainly updates early tokens () for representation abstraction, while T2I gradients flow from later tokens () for detailed reconstruction. Under sequential training (understanding first, generation second), parameter updates are:
with denoting the learning rate. Because both losses propagate through overlapping attention pathways, their gradients are not independent but correlated through shared causal attention weights:
where is the Jacobian of hidden states w.r.t. , and denotes token-level prediction error. When the semantic abstraction (MMU) and detailed reconstruction (T2I) objectives induce opposite attention gradients on overlapping tokens, the inner product
becomes negative, revealing a gradient-level conflict specific to AR architectures, where the resulting degradation in understanding loss after generation updates can be approximated as:
Because under negative inner products, the model exhibits negative transfer and catastrophic forgetting.
3.2. Task-Aware MoE Layer
To better accommodate task objective conflict of understanding and generation tasks, we introduce a modular design based on a MoE framework. Specifically, the expert are divided into two groups, each tailored to understanding and generation tasks, respectively. To enhance task-specific specialization, we propose a Hierarchical Expert Routing mechanism, which comprises a task-aware router and a conventional dynamic assignment router. This mechanism enables more accurate routing of task-relevant tokens to the appropriate expert. Furthermore, to ensure coherence and synergy across task objectives, we incorporate a shared expert independent of the task-specific groups. As a core part of the Hybrid Expert Collaboration mechanism, the shared expert facilitates cross-task information exchange and helps maintain a balanced representation between MMU and T2I tasks.
3.2.1. Hierarchical Expert Routing
To achieve fine-grained expert selection tailored to different task objectives, we introduce Hierarchical Expert Routing, consisting of a Task-Aware Router and a Dynamic-Assignment Router.
Task-Aware Router. Given an input token representation , we first classify it into one of the two task-specific expert groups via a softmax-based classifier:
This essentially acts as a hard routing mechanism, directly determining the group that the token should be routed to.
Dynamic-Assignment Router. Once the expert group is determined, we further compute the relevance between the token and each expert in the group, and select the top- experts based on the scores:
Here, denotes the group-specific expert scoring matrix, where is the number of experts in each group.
3.2.2. Hybrid Expert Collaboration
After selecting experts, it processes input in a weighted combination:
To leverage both task-specific and global knowledge, we integrate the output of a shared expert:
where is a learnable hyperparameter to balance contributions.
Loss Function.
The overall training objective includes both task-specific losses and group loss:
where is the ground truth for the task group, is the loss for task , and , are hyperparameters controlling the loss weights. Specifically, corresponds to the understanding task, while controls the generation task.
3.3. Training Strategy
To fully leverage the unique capabilities of different experts and enhance the generalization ability of the overall framework, we propose a Two-stage Training strategy, as illustrated in Figure 3. In the following paragraphs, we detail objectives and provide a comprehensive explanation of each training stage.
3
Stage 1: Training Task-Specific Experts. This stage concentrates on deriving task-specific experts via specialized training for each individual task. The aim is to obtain expert weights that excel in their corresponding tasks. Mathematically, we employ the generative entropy loss as the main training objective, which can be formulated as:
where denotes the true probability distribution of the i-th sample over C classes, and is the predicted probability from the task-specific expert. During this training phase, the self-attention layers, characterized by their parameter set , are frozen (i.e., ), and only the feed-forward networks (FFNs) with parameter are updated. This leads to the formation of highly specialized task-specific experts, as expressed by the parameter update rule for the FFN:
where is the learning rate and represents the training iteration.
Stage 2: LoRA-Based Fine-Tuning of UniDecouple. This stage, we first replace the traditional FFN layers with the Task-Aware MoE (TAMoE) layer. The TAMoE layer’s output can be described as:
Here, are the weights assigned by the task-aware router to select the top- experts from either the understanding group (Und. Group) or the generation group (Gen. Group), and is the output of the -th expert in group for input . Then, we incorporate the expert weights obtained in Stage 1 and conduct joint fine-tuning of UniDecouple using a combination of understanding and generation tasks.
Specially, we perform end-to-end LoRA fine-tuning with a rank of , enabling UniDecouple to efficiently adapt to diverse tasks with improved generalization through highly parameter-efficient optimization.
4. Experiments
In this section, we present comprehensive experiments for our proposed method. First, we describe the experimental setup, including the model architecture and training configurations. Next, we verify the existence of task objective conflicts (TOC) through metrics such as loss and evaluate the model on various understanding and generation benchmarks. We then provide qualitative results to illustrate the strengths of our approach. Finally, we analyze the utilization of both task-specific and shared experts.
4.1. Experimental Setups
Model Architecture.
We adopt JanusPro as the base model, whose language component supports a maximum sequence length of 4096. For multimodal understanding, we use SigLIP-Large-Patch16-384 Zhai et al. (2023) as the image encoder. For image generation, the encoder utilizes a codebook of size 16,384 and downsamples inputs by a factor of 16. Both understanding and generation adapters are implemented as two-layer MLPs.
Training Configurations.
For the training setup, we keep the vision tower frozen in all stages since it is sufficiently pretrained by Chen et al. (2025). Task-specific experts are trained on ShareGPT4V Chen et al. (2024b) for understanding and MidJourney Vivym (2023) for generation, while the LoRA fine-tuning stage uses a mixed dataset of both. In the MoE design, each task group is assigned two experts with Top- routing, and one shared expert is added to enhance generalization. The hyperparameters are as follows: in Stage 1, batch size of 32, learning rate , and 200 training steps; in Stage 2, batch size of 32, learning rate , and 400 training steps. Both stages use the AdamW optimizer (), cosine learning rate scheduler, and weight decay of 0.0. The main loss coefficients are set to , , and .
4.2. Existence of Conflict
First, we design two variants based on Chen et al. (2025): (1) a single-task model, where each task is trained independently, and (2) a multi-task model, where task-specific losses are jointly optimized. Using POPE and GenEval to cover understanding and generation tasks, results show that the multi-task model suffers performance drops on certain tasks and fails to maintain balanced performance, demonstrating inherent conflicts between task objectives.
| Task | POPE | GenEval | POPE | Gen.Eval | |
| (a) | Und. | 86.9 | – | – | – |
| (b) | Gen. | – | 0.74 | – | – |
| (c) | Both | 86.2 | 0.73 | -0.7 | -0.01 |
Second, to further verify TOC, we analyze the dynamics of the loss functions during multi-task training. Specifically, we track the loss values of multimodal understanding and image generation tasks over training iterations to identify whether decreases in one task’s loss correspond to increases in another’s. Our understanding task uses cross-entropy loss, while the generation task employs mean squared error (MSE) loss. If there is significant mutual suppression in loss curves, it indicates the existence of optimization conflicts between tasks.
From Figure 4, we observe a certain degree of antagonism between the two loss functions, which further supports the existence of task conflicts. However, these experiments do not provide a thorough verification of the conflict, and addressing this at a deeper level remains a goal for our future work.
Loss
4.3. Quantitative Evaluation
Multimodal Understanding (MMU). To assess MMU, we evaluate our model on widely recognized understanding benchmarks, which include POPE Li et al. (2023a), MME Fu et al. (2024), MMB Liu et al. (2024a), SEED Li et al. (2024a), GQA Hudson and Manning (2019), MMMU Yue et al. (2024) and MM-Vet Yu et al. (2023).
Text-to-Image Generation (T2I). For evaluating T2I capabilities, we use GenEval Ghosh et al. (2023) and DPG-Bench Hu et al. (2024).
Comparison Methods. As shown in Table 2 and Table 4, we categorize the models into three groups based on their capabilities. (i) Understanding Only: This group includes models designed solely for MMU, such as the LLaVA series (e.g., LLaVA Liu et al. (2023), LLaVA-v1.5 Liu et al. (2024c)), the MobileVLM series (e.g., MobileVLM Chu et al. (2023), MobileVLM-V2 Chu et al. (2024)), InstructBLIP Dai et al. (2023), Qwen-VL-Chat Bai et al. (2023), and Emu3-Chat Wang et al. (2024c). (ii) Generation Only: This group contains models that focus exclusively on image generation, including LlamaGen Sun et al. (2024), LDM Rombach et al. (2022), SDv1.5 Rombach et al. (2022), PixArt- Chen et al. (2023b), SDv2.1 Rombach et al. (2022), DALL·E 2 Ramesh et al. (2022), Emu3-Gen Wang et al. (2024c), SDXL Podell et al. (2023), DALL·E 3 Betker et al. (2023), and SD3-Medium Esser et al. (2024). (iii) Understanding and Generation: This group comprises models capable of both MMU and T2I, such as DreamLLM Dong et al. (2023), LaViT Jin et al. (2023), MetaMorph Tong et al. (2024), NEXT-GPT Wu et al. (2024b), Show-o-256 Xie et al. (2024), SDHo-512 Xie et al. (2024), D-DiT Li et al. (2024b), Gemini-Nano-1 Team et al. (2023), ILLUME Wang et al. (2024b), TokenFlow-XL Qu et al. (2024), LWM Liu et al. (2024e), VILA-U Wu et al. (2024c), Chameleon Team (2024), and the Janus-series models Wu et al. (2024a); Chen et al. (2025); Ma et al. (2024).
| Type | Model | Params | POPE↑ | MME-P↑ | MMB↑ | SEED↑ | GQA↑ | MMMU↑ | MM-Vet↑ |
| Und. Only | LlaVA-v1.5-Phi-1.5 Xie et al. (2024) | 1.3B | 84.1 | 1128.0 | - | - | 56.5 | 307 | - |
| MobileVLM Chu et al. (2023) | 1.4B | 84.5 | 1196.2 | 53.2 | - | 56.1 | - | - | |
| MobileVLM-V2 Chu et al. (2024) | 1.4B | 84.3 | 1302.8 | 57.7 | - | 59.3 | - | - | |
| MobileVLM Chu et al. (2023) | 2.7B | 84.9 | 1288.9 | 59.6 | - | 59.0 | - | - | |
| MobileVLM-V2 Chu et al. (2024) | 2.7B | 84.7 | 1440.5 | 63.2 | - | 61.1 | - | - | |
| LlaVA-Phi Zhu et al. (2024a) | 2.7B | 85.0 | 1335.1 | 59.8 | - | - | - | 28.9 | |
| LLaVA Liu et al. (2023) | 7B | 76.3 | 809.6 | 38.7 | 33.5 | - | - | 25.5 | |
| LLaVA-v1.5 Liu et al. (2024c) | 7B | 85.9 | 1510.7 | 64.3 | 58.6 | 62.0 | 35.4 | 31.1 | |
| InstructBLIP Dai et al. (2023) | 7B | - | - | 36.0 | 53.4 | 49.2 | - | 26.2 | |
| Qwen-VL-Chat Bai et al. (2023) | 7B | - | 1487.5 | 60.6 | 58.2 | 57.5 | - | - | |
| Emu3-Chat Wang et al. (2024c) | 8B | 85.2 | 1244 | 58.5 | 68.2 | 60.3 | 31.6 | 37.2 | |
| InstructBLIP Dai et al. (2023) | 13B | 78.9 | 1212.8 | - | - | 49.5 | - | 25.6 | |
| Und. and Gen. | DreamLLM† Dong et al. (2023) | 7B | - | - | - | - | - | - | 36.6 |
| LaViT† Jin et al. (2023) | 7B | - | - | - | - | 46.8 | - | - | |
| MetaMorph† Tong et al. (2024) | 8B | - | - | 75.2 | 71.8 | - | - | - | |
| NEXT-GPT† Wu et al. (2024b) | 13B | - | - | - | - | - | - | - | |
| Show-o-256 Xie et al. (2024) | 1.3B | 73.8 | 948.4 | - | - | 48.7 | 25.1 | - | |
| SDHo-512 Xie et al. (2024) | 1.3B | 80.0 | 1097.2 | - | - | 58.0 | 26.7 | - | |
| D-DiT Li et al. (2024b) | 2.0B | 84.0 | 1124.7 | - | - | 59.2 | - | - | |
| Gemini-Nano-1 Team et al. (2023) | 1.8B | - | - | - | - | - | 26.3 | - | |
| TokenFlow-XL Qu et al. (2024) | 13B | 86.8 | 1545.9 | 68.9 | 68.7 | 62.7 | 38.7 | 40.7 | |
| LWM Liu et al. (2024e) | 7B | 75.2 | - | - | 44.8 | - | - | 9.6 | |
| VILA-U Wu et al. (2024c) | 7B | 85.8 | 1401.8 | - | 59.0 | 60.8 | - | 33.5 | |
| Chameleon Team (2024) | 7B | - | - | - | - | - | 22.4 | 8.3 | |
| Janus Wu et al. (2024a) | 1.5B | 87.0 | 1338.0 | 69.4 | 63.7 | 59.1 | 30.5 | 34.3 | |
| JanusPro Chen et al. (2025) | 1.5B | 86.2 | 1444.0 | 75.5 | 68.3 | 59.3 | 36.3 | 39.8 | |
| UniDecouple (Ours) | 3B | 87.2 | 1534.1 | 79.3 | 72.4 | 61.5 | 41.7 | 46.7 |
| Method | Global | Entity | Attribute | Relation | Other | Overall |
| SDv1.5 Rombach et al. (2022) | 74.63 | 74.23 | 75.39 | 73.49 | 67.81 | 63.18 |
| PixArt- Chen et al. (2023b) | 74.97 | 79.32 | 78.60 | 82.57 | 76.96 | 71.11 |
| SDXL Podell et al. (2023) | 83.27 | 82.43 | 80.91 | 86.76 | 80.41 | 74.65 |
| Hunyuan-DiT Li et al. (2024c) | 84.59 | 80.59 | 88.01 | 74.36 | 86.41 | 78.87 |
| PixArt- Chen et al. (2024a) | 86.89 | 82.89 | 88.94 | 86.59 | 87.68 | 80.54 |
| Emu3-Gen Wang et al. (2024c) | 85.21 | 86.68 | 86.84 | 90.22 | 83.15 | 80.60 |
| DALL-E 3 Betker et al. (2023) | 90.97 | 89.61 | 88.39 | 90.58 | 89.83 | 83.50 |
| SD3-Medium Esser et al. (2024) | 87.90 | 91.01 | 88.83 | 80.70 | 88.68 | 84.08 |
| Janus Wu et al. (2024a) | 82.33 | 87.38 | 87.70 | 85.46 | 86.41 | 79.68 |
| JanusPro Chen et al. (2025) | 87.58 | 88.63 | 88.17 | 88.98 | 88.30 | 82.63 |
| UniDecouple (Ours) | 87.95 | 89.07 | 89.31 | 89.33 | 88.84 | 84.11 |
| Type | Method | Single Obj. | Two Obj. | Counting | Colors | Position | Color Attri. | Overall |
| Gen. Only | LlamaGen Sun et al. (2024) | 0.71 | 0.34 | 0.21 | 0.58 | 0.07 | 0.04 | 0.32 |
| LDM Rombach et al. (2022) | 0.92 | 0.29 | 0.23 | 0.70 | 0.02 | 0.05 | 0.37 | |
| SDv1.5 Rombach et al. (2022) | 0.97 | 0.38 | 0.35 | 0.76 | 0.04 | 0.06 | 0.43 | |
| PixArt- Chen et al. (2023b) | 0.98 | 0.50 | 0.44 | 0.80 | 0.08 | 0.07 | 0.48 | |
| SDv2.1 Rombach et al. (2022) | 0.98 | 0.51 | 0.44 | 0.85 | 0.07 | 0.17 | 0.50 | |
| DALL-E2 Ramesh et al. (2022) | 0.94 | 0.66 | 0.49 | 0.77 | 0.10 | 0.19 | 0.52 | |
| Emu3-Gen Wang et al. (2024c) | 0.98 | 0.71 | 0.34 | 0.81 | 0.17 | 0.21 | 0.54 | |
| SDXL Podell et al. (2023) | 0.98 | 0.74 | 0.39 | 0.85 | 0.15 | 0.23 | 0.55 | |
| DALL-E 3 Betker et al. (2023) | 0.96 | 0.87 | 0.47 | 0.83 | 0.43 | 0.45 | 0.67 | |
| SD3-Medium Esser et al. (2024) | 0.99 | 0.94 | 0.72 | 0.89 | 0.33 | 0.60 | 0.74 | |
| Und. and Gen. | SEED-X† Li et al. (2024a) | 0.97 | 0.58 | 0.26 | 0.80 | 0.19 | 0.14 | 0.49 |
| Show-o Xie et al. (2024) | 0.95 | 0.52 | 0.49 | 0.82 | 0.11 | 0.28 | 0.53 | |
| D-DiT Li et al. (2024b) | 0.97 | 0.80 | 0.54 | 0.76 | 0.32 | 0.50 | 0.65 | |
| LWM Liu et al. (2024e) | 0.93 | 0.41 | 0.46 | 0.79 | 0.09 | 0.15 | 0.47 | |
| Transfusion Zhou et al. (2024) | - | - | - | - | - | - | 0.63 | |
| ILLUME Wang et al. (2024b) | 0.99 | 0.86 | 0.45 | 0.71 | 0.39 | 0.28 | 0.61 | |
| TokenFlow-XL Qu et al. (2024) | 0.95 | 0.60 | 0.41 | 0.81 | 0.16 | 0.24 | 0.55 | |
| Chameleon Team (2024) | - | - | - | - | - | - | 0.39 | |
| Janus Wu et al. (2024a) | 0.97 | 0.68 | 0.30 | 0.84 | 0.46 | 0.42 | 0.61 | |
| JanusPro Chen et al. (2025) | 0.98 | 0.82 | 0.51 | 0.89 | 0.65 | 0.56 | 0.73 | |
| UniDecouple (Ours) | 0.99 | 0.84 | 0.55 | 0.91 | 0.62 | 0.65 | 0.76 |
All in all, our UniDecouple, based on Chen et al. (2025), preserves strong MMU ability while achieving T2I quality comparable to state-of-the-art methods. This success is primarily attributed to TA-MoE, which effectively mitigates TOC between MMU and T2I, thereby significantly enhancing model’s overall capabilities.
4.4. Qualitative Evaluation
Figure 6 presents qualitative results for both multimodal understanding (left) and text-to-image generation (right). UniDecouple demonstrates outstanding comprehension capabilities when processing inputs from diverse contexts, showcasing its strong capacity for multimodal semantic modeling, particularly in Basic Visual Description, Visual Question Answering (VQA), and Text Recognition tasks. The right side presents representative T2I results across four categories: basic spatial relationships, natural landscapes, human characters, and imaginative scenes. UniDecouple effectively captures the semantic intent of prompts, generating coherent, well-structured, and detail-rich images. These results highlight the superiority of our approach. UniDecouple not only models high-level semantic information in MMU tasks but also faithfully preserves fine-grained visual details in generation tasks, thereby mitigating inter-TOC.
expert-load
4.5. Expert Load Analysis
| Model | Task-Aware Router | Shared | Understanding | Generation | ||
| + Group Loss | Expert | MMMU | POPE | Gen.eval | ||
| A | ✗ | ✗ | 36.8 | 85.4 | 0.70 | |
| B | ✓ | ✗ | 39.9 | 86.9 | 0.72 | |
| C | ✓ | ✓ | 41.7 | 87.2 | 0.76 | |
| Model | MoE | Train. Method | Expert Init. | Epochs to Convergence | MMMU | POPE | Gen.eval |
| D | ✓ | Single Stage | pure | 94 | 36.5 | 85.3 | 0.68 |
| E | ✓ | Two Stage | from Stage1 | 76 | 41.7 | 87.2 | 0.76 |
Figure 5 illustrates expert utilization in understanding task, visualized by averaging values from randomly selected samples across multiple experimental runs. As shown in 5, Expert-0 and Expert-1 belong to the understanding group, with their combined activation rate exceeding 60% in most cases. Experts 2 and 3, belonging to the generation group, exhibit low activation rates across most layers, indicating that our task-aware router effectively routes tokens to the intended expert groups. Even when occasional routing errors increase generation group activations, the shared expert’s influence can partially compensate for meeting the task requirements. This further validates the effectiveness of our routing mechanism and shared expert design.
5. Ablation Study
In this section, we conduct a comprehensive ablation study to investigate the contributions of key components in our proposed framework. We focus on analyzing the effectiveness of Task-Aware MoE, the impact of training strategy and expert ratio.
5.1. Effectiveness of Task-Aware MoE
Vis
To assess the effectiveness of each component in the Task-Aware MoE, we progressively construct a series of ablation models, as shown in Table 6. We begin with Model A as the baseline, which excludes two critical components: the Task-Aware Router (including Group Loss) and the Shared Expert. Model B introduces the Task-Aware Router, enabling dynamic expert group selection based on task-specific requirements. Building on this, Model C incorporates the Shared Expert module, which promotes cross-task knowledge sharing and enhances the model’s generalization capability.
We evaluate understanding performance using the MMMU and POPE benchmarks, and assess generation quality with the Gen.eval metric, reporting the “overall” score as the primary evaluation criterion. As shown in Table 6, each added component yields consistent performance gains, with Model C achieving the best results across all tasks. This ablation demonstrates the effectiveness and necessity of each module in enhancing both task-specific adaptation and generalization capabilities within the Task-Aware MoE framework.
5.2. Effectiveness of Training Strategy
We investigate the impact of different training strategies on model performance through two distinct model variants presented in Table 6. To ensure a fair comparison, all models share the same architecture and are evaluated on identical task sets. The comparative results between models D and E demonstrate that the two-stage training strategy employed by model E yields significant improvements across various benchmarks. This indicates that incorporating dedicated training stages for handling distinct tasks not only enhances model performance but also accelerates convergence, underscoring the strategic advantage of involving experts in task-specific training.
5.3. Impact of Expert Ratio
| G:S Ratio | POPE | GenEval | |
| (d) | 1:0 | 85.5 | 0.68 |
| (e) | 0:1 | 86.9 | 0.74 |
| (f) | 1:1 | 87.1 | 0.74 |
| (g) | 2:1 | 87.2 | 0.76 |
| (h) | 3:1 | 86.2 | 0.73 |
The visualization in subsection 4.5 illustrates how UniDecouple utilizes experts on the understanding task, yet the optimal balance between group-specific and shared experts has not been fully established. To explore this, we conducted experiments on POPE and GenEval (Table 7), which suggest that a roughly 2:1 ratio yields the best results. Importantly, due to differences in routing strategy, we do not directly adopt the 3:1 ratio reported in DeepSeek Dai et al. (2024). This finding is preliminary and requires further investigation in future work.
6. Conclusion
In this paper, we address the challenge of Task Objective Conflicts in unified multimodal autoregressive models. We propose UniDecouple, a novel framework that decouples internal modules and constructs task-specific optimization subpaths, effectively mitigating negative transfer between multimodal understanding (MMU) and text-to-image generation (T2I) tasks. Central to our approach is the Task-Aware Mixture-of-Experts (TA-MoE), which combines Hierarchical Expert Routing and Hybrid Expert Collaboration, trained with a two-stage strategy to balance task specialization and overall coordination. Extensive experiments demonstrate that UniDecouple preserves strong understanding capabilities while achieving generation quality on par with state-of-the-art methods. Our results highlight the importance of disentangling task-specific pathways in unified models and provide a promising direction for future research on efficient, high-performing multimodal architectures.